OpenStudy (anonymous):

find the prime factorization of 1170 please show all work ??????

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

find the prime factorization of 1170 please show all work ?

OpenStudy (anonymous):

Please

OpenStudy (anonymous):

because the number ends with a zero, you know 1170 is divisible by....

OpenStudy (anonymous):

by what

OpenStudy (anonymous):

Best Answer - Chosen by Voters First step is easy, 10 is a factor, so 1170 = 117 * 10 We know that 10 = 5*2, so 1170 = 117*10 = 117*5*2 117 happens to be divisible by 9 (since the sum of digits is 9). 9 goes in 13 times, so 1170 = 117 * 10 = 117 * 5 * 2 = 9 * 13 * 5 * 2. Of course, 9 is 3* 3, so 1170 = 117 * 10 = 117 * 5 * 2 = 9 * 13 * 5 * 2 = 3 * 3 * 13 * 5 * 2 h

OpenStudy (anonymous):

yay ^^^... by 10...

OpenStudy (anonymous):

wait a min did you just solve it or what? because it says do not solve lol i dont know ......

OpenStudy (anonymous):

yes I solved it you wanted all the steps so I showed you step by step how to factor it

OpenStudy (anonymous):

but if i said 10 in the answer then i would be solving it right ?

OpenStudy (anonymous):

Yes you would solve it but that is how you show your steps. If you have to show steps then you will solve it along the way.

OpenStudy (anonymous):

man ur CEO facebook smart man

OpenStudy (anonymous):

sorry I love math....I am helping my husband at the moment do Algebra so that he can understand it.

OpenStudy (anonymous):

I left you a message

OpenStudy (anonymous):

and only have one more bud and add me to as friend please im new but cool

OpenStudy (anonymous):

I answered for you

OpenStudy (anonymous):

I have no clue how to add you as a friend

OpenStudy (anonymous):

ok i know i just gottit hahaha Thank You for saving me , ohh i only have one more

OpenStudy (anonymous):

I think you can fan me and that way you can always send me questions

OpenStudy (anonymous):

send me the next one

OpenStudy (anonymous):

ok i sent it i hope you understand it lol

OpenStudy (anonymous):

I am not sure what they are asking you to do? Do they want you to solve it or do they have that it equals something?

OpenStudy (anonymous):

It says evaulate 12 _____(but this line is under the 12) 3x+y So , 12 on top and 3x=y on the bottom ?????

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x=y i ment

OpenStudy (anonymous):

3x+y i meant

OpenStudy (anonymous):

on the bottom 3x+ y

OpenStudy (anonymous):

so you want to evaluate